# Distributed memory Tensor-train crossT. Shi, D. Hayes, and J. Qiu

### v1

@@ 1 Introduction
The success and development of computing machines in the past few decades have allowed researchers to deal with large scale and high dimensional data easily. Typically, these data sets are stored as multidimensional arrays called tensors [kolda2009tensor], and a general tensor $\mathcal{X}\in\mathbb{C}^{n_{1}\times\cdots\times n_{d}}$ requires a storage cost of $\prod_{j=1}^{d}n_{j}$ degrees of freedom. This scales exponentially with the dimension $d$, and is often referred to as “the curse of dimensionality”. Therefore, data-sparse tensor formats such as canonical polyadic (CP) [hitchcock1927expression], Tucker [de2000multilinear], tensor-train (TT) [oseledets2011tensor], and tensor networks [evenbly2011tensor] with more complex geometries have been proposed. In particular, the TT format, also known as the matrix product state (MPS) in tensor networks and quantum physics, has a memory footprint that scales linearly with respect to the mode sizes $n_{j}$ and dimension $d$. The TT format is widely used in applications such as molecular simulations [savostyanov2014exact], high-order correlation functions [kressner2015low], partial differential equations [guo2023local], constrained optimization [dolgov2017low, benner2020low], and machine learning [vandereycken2022ttml, novikov2020tensor]. Furthermore, the TT format can be incorporated with extra conditions to form special tensor representations that can capture latent data structures. For example, the quantized TT [dolgov2012fast] format is a combination of the TT format and hierarchical structures, and the tensor chain format  [espig2012note] is a result of alterations on MPS.
@@
where $10\leq\epsilon<$ 1 is an accuracy tolerance [grasedyck2013literature, hackbusch2012tensor] and the bounds hold with stability and high probability. Such algorithms include TT singular value decomposition (TTSVD) [oseledets2011tensor], TT sketching [che2019randomized], and QR or SVD based TT cross [oseledets2010tt, dektor2024collocation]. Another category of TT approximation algorithms is to build $\tilde{\mathcal{X}}$ iteratively with greedy and heuristic approaches, such as TT alternating least squares [holtz2012alternating] and adaptive TT cross [oseledets2010tt, dolgov2020parallel]. Although we do not have theoretical guarantees for convergence or convergence rates, these methods can have good performance in certain scenarios. Particularly, adaptive TT cross is a data-based algorithm, with small complexity cost and especially suitable for sparse giant datasets.
@@
In order to exploit modern computing architectures, researchers have proposed various parallel methods, including both shared memory and distributed memory parallelism, for tensor decomposition in CP [li2017model, smith2015splatt], Tucker and hierarchical Tucker [austin2016parallel, grasedyck2019parallel, ballard2020tuckermpi, kaya2016high], and TT [shi2023parallel, grigori2020parallel, chen2017parallelized, wang2020adtt, dolgov2020parallel] format. In addition, tensor operations in TT format, including addition and multiplication [daas2022parallel], contraction [solomonik2014massively], and recompression [al2023randomized] can be executed in parallel as well. In particular, distributed memory parallelism is an operational vessel for data parallel algorithms, so it is not compatible with tensor algorithms that require all elements at once, such as SVD. The major challenge we address in this paper is to construct an approximation $\tilde{\mathcal{X}}$ in a TT format from large $\mathcal{X}$ with distributed memory parallelism. This allows one to partition $\mathcal{X}$ into smaller blocks so that each process handles a significantly smaller chunk of data. Furthermore, with a successful distributed memory design, all processes can execute individual shared memory parallel algorithms and minimize communications with each other, leading to efficient computational and storage consumption.
@@
In this paper, we propose a new parallel adaptive TT cross algorithm based on a distributed memory framework built upon subtensors. A subtensor is a multilinear generalization of a submatrix, and has been used in the matricized-tensor times Khatri-Rao product (MTTKRP) [ballard2018, ballard2020], hierarchical subtensor decomposition [ehrlacher2021], parallel Tucker decomposition [ballard2020tuckermpi], and parallel TT decomposition [shi2023parallel]. The subtensors can carry out key kernels in adaptive TT cross independently. In the end, results on subtensors are gathered to form the outcome of the entire tensor. For the remainder of this paper, we call this subtensor parallelism, as opposed to dimension parallelism. In fact, one can understand subtensor parallelism as a special type of data parallelism, as the key process is to distribute tensor elements to computing architectures in a regular pattern. Comparatively, dimension parallel algorithms partition computations with respect to the dimensionality of the tensor, and thus the number of processes used actively is limited even in the distributed memory setting. The processes can also encounter severe load imbalance as computations for each dimension may vary significantly. In subtensor parallelism, we construct a multidimensional process grid for subtensor partitioning, which enables us to derive explicit bounds on the bandwidth and communication costs. As a bonus, we can run dimension parallel algorithms in the shared memory setting on each process.
@@
In many applications such as numerical integration in quantum mechanics [meyer1990multi] and inverse problems with uncertainty [stuart2010inverse], and data analysis in statistical science [mccullagh2018tensor] and machine learning [rabanser2017introduction], tensors are often formed without an exact formula and can be extremely sparse. In these cases, researchers develop adaptive TT cross [oseledets2010tt] and dimension parallel adaptive TT cross [dolgov2020parallel] for data-centric TT approximation. Our main contribution in this paper is to develop adaptive matrix and tensor cross approximations within the distributed memory framework using submatrices and subtensors. In particular, we derive two novel communication-efficient iterative procedures to construct matrix cross approximations and show they can recover accurate results. These two iterative procedures are combined with submatrix cross, and are used for TT pivot selection and TT core construction respectively. Especially, we show that local pivots selected on the subtensors still hold the nestedness property to maintain tensor interpolation for all steps and all dimensionalities. Furthermore, we can apply dimension parallel TT core construction on each subtensor in a shared memory setting, and achieve a comparably good approximation for the entire tensor.
@@ 2.1 Tensor notation
We use lower case letters for vectors, capital letters for matrices, and calligraphic capital letters for tensors. For notational simplicity, we use MATLAB-style notation to start index counting from 1, and the symbol “:” in indexing. This includes using $a\!:\!b$ to represent the inclusive index set $\{a,a+1,\ldots,b\}$, and a single “:” to represent all the indices in that dimension from start to end. For example, if $\mathcal{Y}\in\mathbb{R}^{4\times 8\times 10}$, then $\mathcal{Y}(\ :\ ,4\!:\!6,\ :\ )$ or $\mathcal{Y}_{:,4:6,:}$ denotes a subtensor of $\mathcal{Y}$ with size $4\times 3\times 10$. In addition, we use index sets for submatrix and subtensor selection. For example, $A(:,\mathcal{J})$ is a submatrix of $A$ with columns $A_{:,j}$ for all $j\in\mathcal{J}$.
@@ 2.2 Tensor-train format
In other words, each element in $\mathcal{X}$ can be computed as the product of a sequence of matrices. The vector $\boldsymbol{s}=(s_{0},\ldots,s_{d})$ is referred to as the size of the TT cores, and in order for the product to be compatible, we require $s_{0}=s_{d}=1$. This TT representation thus has a storage cost of $\sum_{k=1}^{d}s_{k-1}s_{k}n_{k}$, which is linear with respect to both $d$ and $n_{k}$. In addition, we call a vector $\boldsymbol{r}=(r_{0},\ldots,r_{d})$ the TT rank if $\boldsymbol{r}$ contains entry-by-entry smallest possible values of the TT core size. In practice, the exact TT rank is hard to recover, so we either aim to obtain quasi-optimal TT core size from a given threshold, or build an accurate tensor approximation with small pre-selected TT core size. It is shown in [oseledets2011tensor] that ranks of tensor unfoldings bound the TT rank from above, so we hope to use $\boldsymbol{s}$ that satisfies
@@
[…]
@@ 2.3 Pivot selection and TT cross
A major group of algorithms for TT decomposition is based on singular value decomposition (SVD) and randomized SVD, with deterministic or probabilistic error analysis. In practice, however, these SVD-based algorithms can be extremely expensive and unnecessary. For example, tensors originated from real world datasets are often large and sparse, and we often only want an approximation with an accuracy of a few significant digits. This leads to heuristic data-based methods, such as adaptive cross approximation (ACA). In summary, ACA mainly finds a “skeleton” of original data for approximation. For example, the CUR factorization of a matrix is in the form of
$A\approx CUR=A(:,\mathcal{J})A(\mathcal{I},\mathcal{J})^{-1}A(\mathcal{I},:),$
where $\mathcal{I}$ and $\mathcal{J}$ are two index sets. The key building block in ACA is new pivot selection to iteratively enrich these index sets. Matrix ACA can be extended naturally to adaptive TT cross [oseledets2011tensor]. For conceptual simplicity, all cross approximation algorithms we discuss in this article thereafter are with respect to adaptive cross approximations. There are various metrics for pivot selection, but we focus on the greedy approach outlined in [dolgov2020parallel, Algorithm 2], which aims to find an entry with the largest difference between the current approximation and the actual value. This method finds a quasi-optimal pivot, but is computationally much cheaper than other routines such as max volume selection. Suppose we use $\mathcal{I}_{\leq k}$ and $\mathcal{J}_{>k}$ to denote the selected sets of pivots for the $k$th unfolding of a tensor $\mathcal{X}$ with $11\leq k\leq d-$ 1, and $\mathbb{I}_{\ell}$ to denote the set of all indices for dimension $\ell$, $1\leq\ell\leq d$, then one step of finding new pivots is described in Algorithm 1. Notice that $\mathcal{I}_{\leq 0}$ and $\mathcal{J}_{>d}$ here are empty sets. Additionally, pivots selected this way can be shown to satisfy the nestedness property, which ensures that pivots found in one tensor unfolding can be carried over to subsequent unfoldings. As a result, Algorithm 1 is automatically a dimension parallel algorithm.
@@
1:A tensor $\mathcal{X}\in\mathbb{C}^{n_{1}\times\dots\times n_{d}}$, index sets $(\mathcal{I}_{\leq k},\mathcal{J}_{>k})$ containing current pivots.
2:New pivots $(i^{*}_{\leq k},j^{*}_{>k})$
3:for $11\leq k\leq d-$ 1 do
4:     Use [dolgov2020parallel, Algorithm 2] to find $(i^{*}_{\leq k},j^{*}_{>k})$ on a superblock $X_{k}(\mathcal{I}_{\leq k-1}\times\mathbb{I}_{k},\mathbb{I}_{k+1}\times\mathcal{J}_{>k+1})$, which is seen as a submatrix of the $k$th unfolding of $\mathcal{X}$.
Finally, once all pivots are collected in the index sets $(\mathcal{I}_{\leq k},\mathcal{J}_{>k})$ for $11\leq k\leq d-$ 1, then the tensor approximation $\tilde{\mathcal{X}}$ can be built element-wise as
@@
(3) ×$\times\left[\mathcal{X}(\mathcal{I}_{\leq 2},\mathcal{J}_{>2})\right]^{-1}\cdots\mathcal{X}(\mathcal{I}_{\leq d-1},i_{d}),$
where the element access operator ()()( ) are overloaded for notational simplicity. In this way, the TT cores are
@@
[…]
@@ 3 Subtensor parallelism for TT cross
In this section, we develop subtensor parallel TT cross suitable for the distributed memory framework. Throughout this section, we assume each mode size $n_{j}$ is partitioned evenly into $C_{j}$ pieces for $1\leq j\leq d$. In this way, there are $C=\prod_{j=1}^{d}C_{j}$ subtensors in total with roughly the same size. For the moment, we suppose one process handles one subtensor at a time. This assumption can easily be lifted so that the subtensor grid and the process grid are totally different. In the special case that $d=1$ or $d=2$, these grids are often referred to as 1D or 2D grid respectively in literature [blackford1997scalapack]. In order to clearly describe our distributed memory algorithm, we refer to some MPI terminology and functions for communication patterns. These include

### v2

@@ 1 Introduction
The success and development of computing machines in the past few decades have allowed researchers to deal with high-dimensional data easily. Typically, these data sets are stored as multidimensional arrays called tensors [32], and a general tensor $\mathcal{X}\in\mathbb{C}^{n_{1}\times\cdots\times n_{d}}$ requires a storage cost of $\prod_{j=1}^{d}n_{j}$ degrees of freedom. This scales exponentially with the dimension $d$, and is often referred to as “the curse of dimensionality”. Therefore, data-sparse tensor formats such as canonical polyadic (CP) [28], Tucker [15], tensor-train (TT) [39], and tensor networks [22] with more complex geometries have been proposed. In particular, the TT format, also known as the matrix product state (MPS) in tensor networks and quantum physics, has a memory footprint that scales linearly with respect to the mode sizes $n_{j}$ and the dimension $d$. The TT format is widely used in applications such as molecular simulations [42], high-order correlation functions [33], partial differential equations [26], constrained optimization [18, 8], and machine learning [48, 37]. Furthermore, the TT format can be incorporated with extra conditions to form special tensor representations that can capture latent data structures. For example, the quantized TT [19] format is a combination of the TT format and hierarchical structures, and the tensor chain format  [21] is a result of alterations on MPS.
@@
where $0\leq\epsilon<1$ is an accuracy tolerance [23, 27] and the bounds hold with stability and high probability. Such algorithms include TT singular value decomposition (TTSVD) [39], TT sketching [11], and QR or SVD based TT cross [38, 16]. Another category of TT approximation algorithms is to build $\tilde{\mathcal{X}}$ iteratively with greedy and heuristic approaches, such as TT alternating least squares [29] and adaptive TT cross [38, 17]. Although we do not have theoretical guarantees for convergence or convergence rates, these methods can have good performance in certain scenarios. Particularly, adaptive TT cross is a data-based algorithm, with small complexity cost and especially suitable for sparse, giant datasets.
@@
In order to exploit modern computing architectures, researchers have proposed various parallel methods, including both shared memory and distributed memory parallelism, for tensor decomposition in CP [34, 45], Tucker and hierarchical Tucker [2, 24, 3, 31], and TT [43, 25, 12, 49, 17] format. In addition, tensor operations in TT format, including addition and multiplication [14], contraction [46], and recompression [1] can be executed in parallel as well. In particular, distributed memory parallelism is an operational vessel for data parallel algorithms, so it is not compatible with tensor algorithms that require all elements at once, such as SVD. The major challenge we address in this paper is to construct an approximation $\tilde{\mathcal{X}}$ in a TT format from large $\mathcal{X}$ with distributed memory parallelism. This allows one to partition $\mathcal{X}$ into smaller blocks so that each process handles a significantly smaller chunk of data. Furthermore, with a successful distributed memory design, all processes can execute individual shared memory parallel algorithms and minimize communications with each other, leading to efficient computational and storage consumption.
@@
In this paper, we propose a new parallel adaptive TT cross algorithm based on a distributed memory framework built upon subtensors. This is viewed as a direct parallelization of the construction in [38, Section 3]. A subtensor is a multilinear generalization of a submatrix, and has been used in the matricized-tensor times Khatri-Rao product (MTTKRP) [4, 5], hierarchical subtensor decomposition [20], parallel Tucker decomposition [3], and parallel TT decomposition [43]. The subtensors can carry out key kernels in adaptive TT cross independently. In the end, results on subtensors are gathered to form the outcome of the entire tensor. For the remainder of this paper, we call this subtensor parallelism, as opposed to dimension parallelism. In fact, one can understand subtensor parallelism as a special type of data parallelism, as the key process is to distribute tensor elements to computing architectures in a regular pattern. Comparatively, dimension parallel algorithms partition computations with respect to the dimensionality of the tensor, and thus the number of processes used actively is limited even in the distributed memory setting. The processes can also encounter severe load imbalance as computations for each dimension may vary significantly. In subtensor parallelism, we construct a multidimensional process grid for subtensor partitioning, which enables us to derive explicit bounds on the bandwidth and communication costs. As a bonus, we can run dimension parallel algorithms in the shared memory setting on each process.
@@
In many applications such as numerical integration in quantum mechanics [36] and inverse problems with uncertainty [47], and data analysis in statistical science [35] and machine learning [40], tensors are often formed without an exact formula and can be extremely sparse. In these cases, researchers develop adaptive TT cross [38] and dimension parallel adaptive TT cross [17] for data-centric TT approximation. Our main contribution in this paper is to develop adaptive matrix and tensor cross approximations within the distributed memory framework using submatrices and subtensors. In particular, we derive two novel communication-efficient iterative procedures to construct matrix cross approximations and show they can recover accurate results. These two iterative procedures are combined with submatrix cross, and are used for TT pivot selection and TT core construction respectively. Especially, we show that local pivots selected on the subtensors still hold the nestedness property to maintain tensor interpolation for all steps and all dimensionalities. Furthermore, we can apply dimension parallel TT core construction on each subtensor in a shared memory setting, and achieve a comparably good approximation for the entire tensor.
@@ 2.1 Tensor notation
We use lower case letters for vectors, capital letters for matrices, and calligraphic capital letters for tensors. Just as tensors are higher order analogues of matrices, subtensors are also higher order analogues of submatrices, storing parts of an entire tensor. For notational simplicity, we use MATLAB-style notation to start index counting from 1, and the symbol “:” in indexing. This includes using a:ba\!:\!b to represent the inclusive index set $\{a,a+1,\ldots,b\}$, and a single “:” to represent all the indices in that dimension from start to end. For example, if $\mathcal{Y}\in\mathbb{R}^{4\times 8\times 10}$, then $\mathcal{Y}(\ :\ ,4\!:\!6,\ :\ )$ or $\mathcal{Y}_{:,4:6,:}$ denotes a subtensor of $\mathcal{Y}$ with size $4\times 3\times 10$. In addition, we use index sets for submatrix and subtensor selection. For example, $A(:,\mathcal{J})$ is a submatrix of $A$ with columns $A_{:,j}$ for all $j\in\mathcal{J}$.
@@
Table 1: Notation conventions for tensors and index sets.
Concepts Notations
Vector Lower case letter $u,v,s$
Matrix Capital letter $A,C,U,R$
Tensor Calligraphic capital letter $\mathcal{X},\mathcal{G},\mathcal{T}$
Single index Lower case letter $i,j,k,\ell$
Index set Calligraphic capital letter $\mathcal{I},\mathcal{J},\mathcal{K},\mathcal{L}$
Index set with all possible indices Blackboard bold capital letter $\mathbb{I},\mathbb{J}$
$k$-th Tensor unfolding Capital letter with subscript $X_{k}$
@@ 2.2 Tensor-train format
In other words, each element in $\mathcal{X}$ can be computed as the product of a sequence of matrices. The vector $\boldsymbol{s}=(s_{0},\ldots,s_{d})$ is referred to as the size of the TT cores, and in order for the product to be compatible, we require $s_{0}=s_{d}=1$. This TT representation thus has a storage cost of $\sum_{k=1}^{d}s_{k-1}s_{k}n_{k}$, which is linear with respect to both $d$ and $n_{k}$. In addition, we call a vector $\boldsymbol{r}=(r_{0},\ldots,r_{d})$ the TT rank if $\boldsymbol{r}$ contains entry-by-entry smallest possible values of the TT core size [39, 44]. In practice, the exact TT rank is hard to recover, so we either aim to obtain quasi-optimal TT core size from a given threshold, or build an accurate tensor approximation with small pre-selected TT core size. It is shown in [39] that ranks of tensor unfoldings bound the TT rank from above, so we hope to use $\boldsymbol{s}$ that satisfies
@@
$\mathcal{X}_{i_{1},\ldots,i_{d}}$$\mathcal{G}_{1}(i_{1},:)$$s_{0}\!\times\!s_{1}$$\mathcal{G}_{2}(:,i_{2},:)$$s_{1}\!\times\!s_{2}\cdots$$\mathcal{G}_{d-1}(:,i_{d-1},:)$$s_{d-2}\!\times\!s_{d-1}$
@@ 2.3 Pivot selection and TT cross
A major group of algorithms for TT decomposition is based on singular value decomposition (SVD) and randomized SVD, with deterministic or probabilistic error analysis. In practice, however, these SVD-based algorithms can be extremely expensive and unnecessary. For example, tensors originated from real world datasets are often large and sparse, and we often only want an approximation with an accuracy of a few significant digits. This leads to heuristic data-based methods, such as adaptive cross approximation (ACA). In summary, ACA mainly finds a “skeleton” of original data for approximation. For example, using ACA, one can build the CUR factorization of a matrix, which is in the form of
(3) $A\approx CUR=A(:,\mathcal{J})A(\mathcal{I},\mathcal{J})^{-1}A(\mathcal{I},:),$
where $\mathcal{I}$ and $\mathcal{J}$ are two index sets. The key building block in ACA is new “pivot” selection to iteratively enrich these index sets. The selected indices to represent the data in ACA are referred to as “pivots” due to their connection with the LU factorization. In fact, as shown later in the paragraph, the greedy approach to obtain “pivots” is the same as performing multiple rounds of partial pivoting in Gaussian elimination. Matrix ACA can be extended naturally to adaptive TT cross [38]. For conceptual simplicity, all cross approximation algorithms we discuss in this article thereafter are with respect to adaptive cross approximations. There are various metrics for pivot selection, but we focus on the greedy approach outlined in [17, Algorithm 2] (we include as Algorithm 1 for self-containment), which aims to find an entry with the largest difference between the current approximation and the actual value. This method finds a quasi-optimal pivot through heuristics, and thus there is not a theoretical guarantee for accuracy, but is computationally much cheaper than other routines such as max volume selection. Suppose we use $\mathcal{I}_{\leq k}$ and $\mathcal{J}_{>k}$ to denote the selected sets of pivots for the row and column indices of $X_{k}$, the $k$th unfolding of a tensor $\mathcal{X}$ with $1\leq k\leq d-1$, and $\mathbb{I}_{\ell}$ to denote the set of all indices for dimension $\ell$ with $1\leq\ell\leq d$, then one step of finding new pivots is described in Algorithm 2. Notice that $\mathcal{I}_{\leq 0}$ and $\mathcal{J}_{>d}$ here are empty sets, entries of $\mathcal{I}_{\leq k}$ fall between 1 and $\prod_{j=1}^{k}n_{j}$, and entries of $\mathcal{J}_{>k}$ fall between 1 and $\prod_{j=k+1}^{d}n_{j}$. For a pictorial illustration of the keyword “superblock” used in Algorithm 2, we refer the readers to [17]. Additionally, pivots selected this way can be shown to satisfy the nestedness property, which ensures that pivots found in one tensor unfolding can be carried over to subsequent unfoldings. As a result, Algorithm 2 is automatically a dimension parallel algorithm.
Algorithm 1 One step of finding new pivots in matrix cross.
1:A matrix $X\in\mathbb{C}^{n_{1}\times n_{2}}$, index sets $(\mathcal{I},\mathcal{J})$ containing current pivots, and current approximation $\tilde{X}$
[…]
3:Pick a random set of samples $\mathcal{L}$ and select $(i^{*},j^{*})\leftarrow\operatorname*{arg\,max}_{(i,j)\in\mathcal{L}}|X(i,j)-\tilde{X}(i,j)|$.
4:while $|X(i^{*},j^{*})-\tilde{X}(i^{*},j^{*})|<|X(i,j)-\tilde{X}(i,j)|$ for all $(i,j)$ such that $i=i^{*}$ or $j=j^{*}$ do
5:  Select $(i^{*},j^{*})\leftarrow\operatorname*{arg\,max}_{i\in\mathbb{I}}|X(i,j^{*})-\tilde{X}(i,j^{*})|$.
6:  Select $(i^{*},j^{*})\leftarrow\operatorname*{arg\,max}_{j\in\mathbb{J}}|X(i^{*},j)-\tilde{X}(i^{*},j)|$.
@@
1:A tensor $\mathcal{X}\in\mathbb{C}^{n_{1}\times\dots\times n_{d}}$, index sets $(\mathcal{I}_{\leq k},\mathcal{J}_{>k})$ containing current pivots for $1\leq k\leq d-1$.
2:New pivots $(i^{*}_{\leq k},j^{*}_{>k})$ for $1\leq k\leq d-1$.
3:for $1\leq k\leq d-1$ do
4:  Use Algorithm 1 to find $(i^{*}_{\leq k},j^{*}_{>k})$ on a superblock $X_{k}(\mathcal{I}_{\leq k-1}\times\mathbb{I}_{k},\mathbb{I}_{k+1}\times\mathcal{J}_{>k+1})$, which is seen as a submatrix of the $k$th unfolding of $\mathcal{X}$.
Finally, once all pivots are collected in the index sets $(\mathcal{I}_{\leq k},\mathcal{J}_{>k})$ for $1\leq k\leq d-1$, then the tensor approximation $\tilde{\mathcal{X}}$ can be built element-wise as
@@
(4) $\times\left[\mathcal{X}(\mathcal{I}_{\leq 2},\mathcal{J}_{>2})\right]^{-1}\cdots\mathcal{X}(\mathcal{I}_{\leq d-1},i_{d}),$
where the element access operator $()$ are overloaded for notational simplicity, and all inverses are matrix inverses. The use of overloading allows for consistent notation to avoid mixing terms to select 3-way tensors with $\mathcal{X}(\cdot,\cdot,\cdot)$ and matrices with $X_{k}(\cdot,\cdot)$. This can be seen as an extension of Eq. 3 in the TT format, with multiple subtensor selections and cross section inverses. In this way, we can group consecutive terms in the previous expression to construct the TT cores as
@@
(5) $\mathcal{G}_{k}=$ $\mathcal{X}(\mathcal{I}_{\leq k-1},:,\mathcal{J}_{>k})\left[\mathcal{X}(\mathcal{I}_{\leq k},\mathcal{J}_{>k})\right]^{-1},\ 2\leq k\leq d-1.$
@@ 3 Subtensor parallelism for TT cross
In this section, we develop subtensor parallel TT cross suitable for the distributed memory framework. In particular, the distributed computing processes handle non-overlapping subtensors, but work together to obtain the TT format of the entire tensor. Throughout this section, we assume each mode size $n_{j}$ is partitioned evenly into $C_{j}$ pieces for $1\leq j\leq d$. In this way, there are $C=\prod_{j=1}^{d}C_{j}$ subtensors in total with roughly the same size. For the moment, we suppose one process handles one subtensor at a time. This assumption can easily be lifted so that the subtensor grid and the process grid are totally different. In the special case that $d=1$ or $d=2$, these grids are often referred to as 1D or 2D grid respectively in literature [9]. In order to clearly describe our distributed memory algorithm, we refer to some MPI terminology and functions for communication patterns. These include